OpenStudy (anonymous):

4(2-6m)=6(2-4m) show work

OpenStudy (chaise):

4(2-6m)=6(2-4m) Distrubute your brackets. 8-24m=12-24m Plus 24m both sides 8=12 Therefore fale; no solutions exist.

OpenStudy (anonymous):

Please show steps

OpenStudy (precal):

\[8\neq12\]\[8-24m=12-24m\]\[4(2)-4(6m)=6(2)-6(4m)\]\[4(2-6m)=6(2-4M)\] Note: There is no value of m that will ever make these two expressions equal each other. Try subbing any values and you will see there is no possible soltuion. By the way Chaise did a great job of explaining.